OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –6 –6 –4 –3 Line 2 x y 0 3 2 6 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (solomonzelman):

can you tell me the slope of the second line

OpenStudy (anonymous):

3/2

OpenStudy (solomonzelman):

slope: line 1 (-6 - -3) /-6 -- 4) >> -3/-2 >> 3/2 line 2: (6 - 3) / (2 - 0) >> 3/2

OpenStudy (solomonzelman):

both lines have a slope of -3/2

OpenStudy (solomonzelman):

ops, just 3/2

OpenStudy (solomonzelman):

find the equartion of the first line using: \(\large\color{slate}{ y-y_1=m(x-x_1)}\)

OpenStudy (anonymous):

Um. What?

OpenStudy (solomonzelman):

x y –6 –6 –4 –3 \(\normalsize\color{slate}{ \rm y-(-4)=\frac{3}{2}(x-(-3)) }\) \(\normalsize\color{slate}{ \rm y+4=\frac{3}{2}(x+3) }\) \(\normalsize\color{slate}{ \rm y+4=\frac{3}{2}x+2 }\) \(\normalsize\color{slate}{ \rm y=\frac{3}{2}x-2 }\)

OpenStudy (solomonzelman):

refresh to get rid of question marks.

OpenStudy (solomonzelman):

so the y intercepts of these 2 paralle lines are different, thus they are not on top of each other

OpenStudy (solomonzelman):

that means the lines won't intersecrt

OpenStudy (solomonzelman):

intersect*

OpenStudy (anonymous):

So.....?

OpenStudy (solomonzelman):

a solution is an intersection of the 2 graphs

OpenStudy (anonymous):

So that means no solution?